# Different physical and numerical sources of scatter in the $M_{\star}$-$M_{\mathrm{BH}}$ relation and their connection to galaxy evolution

### v1

@@ 2.2.1 Data selection
For the sample taken from MTNG based on the selection criteria above we find that there is a comparatively prominent subpopulation with lower BH mass compared to the “main sequence” of the $M_{\rm BH}-M_{\rm\star}$ relation, as shown in the upper panels of Figure 1. Most of these galaxies belong to massive galaxy clusters and are not central galaxies, and further investigation reveals that they have a very similar color compared to the galaxies with similar stellar mass on the main sequence. While a deep investigation of their formation history is beyond the scope of the current work, we believe that the much smaller abundance of these systems in TNG is related to a known deficiency of its black hole reposition algorithm (Borrow et al., 2023a), which can make satellite galaxies in clusters lose their black hole prematurely once all of their gas is stripped, a problem that has been fixed in MTNG.
For simplicity, we here use a Gaussian mixture model (GMM) clustering algorithm to separate these satellite galaxies from the samples taken from MTNG. The bottom panel of Figure 1 shows that the GMM performs well for this task, and so we use the resulting cleaned MTNG-sample for our analyses below.
@@ 2.2.2 Fitting procedure
The data samples and the fitting results from $z=3$ to $z=0$ in TNG, Illustris, EAGLE and MTNG with the selection criteria $\sigma>150\,{\rm km\,s^{-1}}$ and $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$ are presented in Figure 2. For the fitting procedure, we adopt a linear model in the log as in Kormendy & Ho (2013), in the form
@@
Similar to observational data, which has individual measurement errors, the simulation data has also errors that can be identified with the uncertainties originating in numerical variations or stochastic modeling, i.e. $M_{\rm BH}+\delta_{\rm BH}$ and $M_{\star}+\delta_{\star}$, where $\delta_{\rm BH}$ and $\delta_{\star}$ are variations introduced by the numerical discretization scheme or the use of stochastic sub-grid modeling. Due to the existence of these errors, one should in principle not use the simulation data with the simple linear regression method as this can yield a biased estimated of the slope. Indeed, if the simple linear regression is used, the estimated slope $\hat{\beta}$ becomes
@@
Since we do not have any reliable information about the size of the errors due to the numerical variations and stochastic modeling, we utilize a grouping method to obtain an unbiased estimate of slope. The procedure of the grouping method is as follows:
@@
  • Apply a weighted least square method (WLS) to do a linear regression for the bin-averaged data.
Since the errors from numerical variations and stochastic modeling become zero when averaging the data within each bin, the simple linear regression for the bin-averaged data should yield an unbiased estimate of the slope.
To confirm the effectiveness of this method, we use it to fit the observational data of massive galaxies and classical bulges reported in Kormendy & Ho (2013) and compare with other methods. We use the bisector method (Isobe et al., 1990), $\chi^{2}$ estimator method (Tremaine et al., 2002), and grouping method introduced here to determine the slope and intrinsic scatter, with results shown in Table 1.
@@
Figure 3: Upper panel: The time evolution of scatter in the $M_{\mathrm{BH}}-M_{\star}$ relation in the TNG100, Illustris, EAGLE, and MTNG simulations, obtained by using the fitting procedure mentioned in Section 2.2.2. Lower panel: The time evolution of the corresponding slope in TNG100, Illustris, EAGLE, and MTNG. The solid lines represent the results obtained from the data samples with $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$, while the dashed-dotted lines give the results obtained from data samples with the cut $M_{\star}>10^{11}\,{\rm M_{\odot}}$.
@@ 2.3.1 Time evolution
The fitting results for the data with the $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$ selection criterion at different redshifts and for different stellar mass cuts are also included in Figure 2. For the samples with $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$, it can be seen from the figures that the four simulations can be roughly divided into two groups. TNG100 and MTNG have similar values for scatter and slope, with the scatter in both simulations varying from $\sim$0.25 to $\sim$0.15 dex, and the slopes ranging from $\sim$1 to $\sim$0.75 between $z=3$ and $z=0$. In comparison, EAGLE and Illustris show higher scatter, and also a somewhat steeper slope. The values of the scatter in EAGLE vary from $\sim 0.6$ to $\sim$0.3 dex from $z=3$ to $z=0$. In Illustris, the value of the scatter increases from $\sim 0.3$ to $\sim 0.5$ dex over $z=3$ to $z=1$, but then decreases from $\sim 0.5$ to $\sim 0.35$ dex over the range $z=1$ to $z=0$. With respect to the evolution of the slope, the values in Illustris and EAGLE tend to be higher than unity at high redshift, and then decrease to $\sim 1$ by $z=0$.
Since the TNG100 and MTNG simulations share a very similar galaxy formation model with only the box size and the numerical resolution being different, the similar time evolution of scatter and slope indicates a good consistency of the models for the predicted time evolution of the $M_{\rm BH}-M_{\star}$ relation, which is also in line with the findings in Pakmor et al. (2023). Since galaxy mergers are expected to reduce the scatter and make the slope closer to 1, the results for TNG100 and MTNG indicate that mergers mainly contribute to the growth of BHs only at the massive end in TNG100 and MTNG (Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011). The Illustris and EAGLE simulations on the other hand show a larger scatter and steeper slope compared to TNG and MTNG, suggesting that star formation or BH accretion still significantly contribute to the growth of the massive galaxies in these simulations, or in other words, that they are less strongly quenched than corresponding galaxies in TNG and MTNG.
Although we expect that the scatter is not influenced by our estimation procedure of the slope, the intrinsic scatter is still hard to obtain from the simulation data since it may also have contributions from numerical variations and stochastic modeling, described by the last three terms on the right hand side of equation (6). To compare with the intrinsic scatter of observations, we thus need to estimate the numerical contributions to the scatter. We will therefore try to separate the contributions to the scatter coming from different sources, and present a detailed analysis of this point in Section 3.
@@ 3.1 Possible sources of scatter
In equation (6), the second and third terms on the right hand side disappear if the estimation of the slope is unbiased, then only two terms are left. The last term adds scatter due to numerical variations and stochastic modeling, which can be classified as being of purely numerical origin. The numerical sensitivities arise due to the highly non-linear nature of the gravity-hydrodynamical processes in galaxy formation. Here small variations originating in minor differences from numerical round-off or force errors from the TreePM gravity solver can be amplified significantly, which is often referred to as chaotic behaviour or butterfly effects (Genel et al., 2019; Keller et al., 2019; Borrow et al., 2023b). Note also that due to computational limitations, it is impossible to model many astrophysical processes like star formation or BH accretion from first principles. Instead, these phenomena occur ‘sub-grid’, and are often modeled with the help of stochastic treatments. But introducing pseudo random numbers in cosmological simulations induces additional variations in the physical quantities inferred from the simulations, because individual resolution elements do not get to see the same random numbers any more when computations have already started to slightly diverge due to the more subtle reasons mentioned earlier.
@@
The first term in equation (6) represents the intrinsic scatter that comes from the astrophysical processes themselves. The origin of this intrinsic scatter can be broadly divided into two main processes – hierarchical merging on one side, and BH accretion on the other side. The galaxies and their central BHs experience a series of mergers across the cosmological evolution. Since the galaxies do not strictly lie on the scaling relation when they merge, the merger remnant realizes scatter for the galaxies. On the other hand, although the existence of a co-evolution of BHs and their host galaxies may regulate galaxy evolution, making galaxies evolve towards the $M_{\rm BH}-M_{\star}$ relation and reducing the scatter, the BH accretion in itself can also evolve BHs away from the relation and introduce scatter in the scaling relation. The degree to which this happens depends on the physics model for galaxy evolution, and in particular on the feedback for BH growth and star formation. Aside from these physical processes, another important source for scatter in the $M_{\rm BH}-M_{\star}$ relation can lie in variations of the BH seed mass, especially when these seeds are comparatively massive so that substantial galaxy growth is needed before all memory of the seed variations is lost.
@@ 3.2 Numerical experiments
To better characterize different sources of scatter in the $M_{\rm BH}-M_{\star}$ relation, we carry out different numerical experiments, using the TNG model as a fiducial test case. We begin by estimating the scatter from numerical variations and stochastic modeling. To this end we run five TNG box simulations with a box size of $25\,h^{-1}{\rm cMpc}$, using identical initial condition but different numbers of compute nodes. The latter is a means to introduce implicit numerical variations because the AREPO code is only binary invariant in its results if the number of compute nodes/cores is kept unchanged for a repeated simulation. Varying this will change the domain decomposition, and thus the exact order in which gravitational forces and hydrodynamical fluxes are added up, both of which lead to floating point round-off variations. In addition, the random numbers that will be applied to individual gaseous resolution elements for the stochastic modeling of star formation will be completely different when the number of cores for parallel execution is changed.
For the five results we then perform an object-by-object subhalo matching and calculate the variance and standard deviation for the sample of matched subhalos. Since the initial conditions are identical, the variation in galaxy properties for the matched subhalos can be attributed to numerical variations and stochastic modeling. Note that Genel et al. (2019) reported that the variations in galaxy properties measured in TNG box simulations through a similar method is not sensitive to the employed numerical resolution.
@@
The left panel of Figure 5 shows the standard deviation of the BH mass in our five TNG simulation with identical initial condition, as mentioned above. The galaxies are also fulfilling the selection criteria mentioned in Section 2.2.1. We find that the variations of the BH mass roughly follow a Gaussian distribution, except that there are also some data points far away from zero. Including these data points would overestimate the variance of the BH mass, so we remove outliers outside a $3\sigma$ region. As shown in the right panel of Figure 5, the distribution of the data is well preserved when these outliers are removed. Figure 6 compares the time evolution of the measured variance of the BH mass before and after cleaning of the data as mentioned above. When outliers are removed, the variance shows a well behaved behaviour with a value that is roughly constant across $z=3$ to $z=0$, which is consistent with the findings in Genel et al. (2019). We will thus adopt this data cleaning strategy for estimating the size of the scatter due to numerical noise, denoted $\sigma^{2}_{\rm BH}$ in equation (6). Subtracting this from the measured full scatter $\hat{\epsilon}^{2}$, we can then also obtain an estimate of the intrinsic physical scatter $\epsilon^{2}$, which can be viewed as being composed of two main sources, hierarchical merging and gas accretion.
@@
We have additionally run five further TNG box simulations using a similar treatment as mentioned above, but turning the BH accretion off. For these simulations, variations due to BH accretion are not introduced by construction, such that the scatter in these simulations is only caused by hierarchical merging and numerical variations. Assuming that the scatter from different sources is approximately uncorrelated and additive, we can estimate the scatter just from BH accretion by subtracting the corresponding result from the full scatter obtained earlier. Here we for simplicity ignore the stellar mass dependence of the scatter, so this is evidently only a course estimate. But using this approach we can separately estimate the contributions to the intrinsic scatter from hierarchical merging and gas accretion.
@@
Figure 6: The time evolution of the variance of the BH mass in matched halos of five identical TNG simulation with and without data cleaning.

### v2

@@ 2.2.1 Data selection
Throughout this work, we define the BH mass as the sum of the subgrid masses of all BH particles within a given subhalo. In practice, this quantity is typically dominated by a single central BH, so the difference between summing all BH particles and considering only the most massive one is negligible.
For the sample taken from MTNG based on the selection criteria above we find that there is a comparatively prominent subpopulation with lower BH mass compared to the “main sequence” of the $M_{\rm BH}-M_{\rm\star}$ relation, as shown in the upper panels of Figure 1. Most of these galaxies belong to massive galaxy clusters and are not central galaxies, and further investigation reveals that they have a very similar color compared to the galaxies with similar stellar mass on the main sequence. While a deep investigation of their formation history is beyond the scope of the current work, we believe that the much smaller abundance of these systems in TNG100 is related to a known deficiency of its black hole reposition algorithm (Borrow et al., 2023a), which can make satellite galaxies in clusters lose their black hole prematurely once all of their gas is stripped, a problem that has been fixed in MTNG.
For simplicity, we here use a Gaussian mixture model (GMM) clustering algorithm to separate these satellite galaxies from the samples taken from MTNG. The bottom panel of Figure 1 shows that the GMM performs well for this task, and so we use the resulting cleaned MTNG-sample for our analyses below. This selection is motivated by observational limitations: previous observational samples, such as those compiled in Kormendy & Ho (2013), were strongly biased against the detection of undermassive black holes in massive host galaxies, particularly satellite systems. For consistency with these observational constraints, we restrict our comparison sample in MTNG to galaxies near the observed $M_{\rm BH}-M_{\star}$ main sequence.
@@ 2.2.2 Fitting procedure
The data samples and the fitting results from $z=3$ to $z=0$ in TNG, Illustris, EAGLE and MTNG with the selection criteria $\sigma>150\,{\rm km\,s^{-1}}$ and $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$ are presented in Figure 2. For the fitting procedure, we adopt a linear relationship between the logarithms of the masses as in Kormendy & Ho (2013), in the form
@@
Similar to observational data, which have individual measurement errors, the simulation data has also errors that can be identified with the uncertainties originating in numerical variations or stochastic modeling, i.e. $M_{\rm BH}+\delta_{\rm BH}$ and $M_{\star}+\delta_{\star}$, where $\delta_{\rm BH}$ and $\delta_{\star}$ are variations introduced by the numerical discretization scheme or the use of stochastic subgrid modeling. In this context, these variations are analyzed in logarithmic space, i.e., we consider variations in $\log M_{\rm BH}$ and $\log M_{\star}$. In the rest of this paper, all scatter/variance values refer to the logarithmic scatter, defined as the standard deviation in $\log M_{\rm BH}$ at fixed $\log M_{\star}$. Due to the existence of these errors, one should in principle not use the simulation data with the simple linear regression method as this can yield a biased estimate of the slope. Indeed, if the simple linear regression is used, the estimated slope $\hat{\beta}$ becomes
@@
Since we do not have any reliable information about the size of the errors due to the numerical variations and stochastic modelling, we utilize a grouping method to obtain an unbiased estimate of slope. The procedure of the grouping method is as follows:
@@
  • Apply a weighted least-square method (WLS) to do a linear regression for the bin-averaged data.
Since the errors from numerical variations and stochastic modelling become zero when averaging the data within each bin, the simple linear regression for the bin-averaged data should yield an unbiased estimate of the slope.
To confirm the effectiveness of this method, we use it to fit the observational data of massive galaxies and classical bulges reported in Kormendy & Ho (2013) and compare with other methods. We use the bisector method (Isobe et al., 1990), $\chi^{2}$ estimator method (Tremaine et al., 2002), and the grouping method introduced here to determine the slope and intrinsic scatter, with results shown in Table 1. Since our simulated data lacks explicit measurement errors at this stage, this comparison is only performed on the observational dataset.
@@
Figure 3: Upper panel: The time evolution of scatter in the $M_{\mathrm{BH}}-M_{\star}$ relation in the TNG100, Illustris, EAGLE, and MTNG simulations, obtained by using the fitting procedure mentioned in Section 2.2.2. Lower panel: The time evolution of the corresponding slope in TNG100, Illustris, EAGLE, and MTNG. The solid lines represent the results obtained from the data samples with $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$, while the dashed-dotted lines give the results obtained from data samples with the cut $M_{\star}>10^{11}\,{\rm M_{\odot}}$. These results show that the scatter and slope of the $M_{\mathrm{BH}}-M_{\star}$ relation in four simulations can be clearly separated into two groups, independent of the selection criteria.
@@ 2.3.1 Time evolution
The fitting results for the data with the $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$ selection criterion at different redshifts and for different stellar mass cuts are also included in Figure 2. For the samples with $M_{\star}>10^{10.5}\,{\rm M_{\odot}}$, it can be seen from the figures that the four simulations can be roughly divided into two groups. TNG100 and MTNG have similar values for scatter and slope, with the scatter in both simulations varying from $\sim$0.25 to $\sim$0.15 dex, and the slopes ranging from $\sim$1 to $\sim$0.75 between $z=3$ and $z=0$. Despite the significant resolution difference between the two simulations, this consistency suggests that the slope and scatter of the $M_{\rm BH}-M_{\star}$ relation are robust to resolution effects and reflect more fundamental aspects of BH–galaxy co-evolution. In comparison, EAGLE and Illustris show higher scatter, and also a somewhat steeper slope. The values of the scatter in EAGLE vary from $\sim 0.6$ to $\sim$0.3 dex from $z=3$ to $z=0$. In Illustris, the value of the scatter increases from $\sim 0.3$ to $\sim 0.5$ dex over $z=3$ to $z=1$, and then decreases from $\sim 0.5$ to $\sim 0.35$ dex over the range $z=1$ to $z=0$, although the latter is not statistically significant based on a t-test. With respect to the evolution of the slope, the values in Illustris and EAGLE tend to be higher than unity at high redshift, and then decrease to $\sim 1$ by $z=0$.
Since the TNG100 and MTNG simulations share a very similar galaxy formation model with only the box size and the numerical resolution being different, the similar time evolution of scatter and slope indicates a good consistency of the models for the predicted time evolution of the $M_{\rm BH}-M_{\star}$ relation, which is also in line with the findings in Pakmor et al. (2023). Since galaxy mergers are expected to reduce the scatter and make the slope closer to 1 (Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011), the results for TNG100 and MTNG indicate that mergers mainly contribute to the growth of BHs only at the massive end in TNG100 and MTNG. The Illustris and EAGLE simulations on the other hand show a larger scatter and steeper slope compared to TNG100 and MTNG, suggesting that star formation or BH accretion still significantly contribute to the growth of the massive galaxies in these simulations, or in other words, that they are less strongly quenched than corresponding galaxies in TNG100 and MTNG. (See also the discussion in Section 4 and Figure 11.)
Although we expect that the scatter is not influenced by our estimation procedure of the slope, the intrinsic scatter is still hard to obtain from the simulation data since it may also have contributions from numerical variations and stochastic modelling, described by the last three terms on the right hand side of equation (6). To compare with the intrinsic scatter of observations, we thus need to estimate the numerical contributions to the scatter. We will therefore try to separate the contributions to the scatter coming from different sources, and present a detailed analysis of this point in Section 3.
@@ 3.1 Possible sources of scatter
In equation (6), the second and third terms on the right hand side disappear if the estimation of the slope is unbiased, then only two terms are left. The last term adds scatter due to numerical variations and stochastic modelling, which can be classified as being of purely numerical origin. The numerical sensitivities arise due to the highly non-linear nature of the gravity-hydrodynamical processes in galaxy formation. Here small variations originating in minor differences from numerical round-off or force errors from the TreePM gravity solver can be amplified significantly, which is often referred to as chaotic behaviour or butterfly effects (Genel et al., 2019; Keller et al., 2019; Borrow et al., 2023b). Note also that due to computational limitations, it is impossible to model many astrophysical processes like star formation or BH accretion from first principles. Instead, these phenomena occur “subgrid”, and are often modeled with the help of stochastic treatments. But introducing pseudo random numbers in cosmological simulations induces additional variations in the physical quantities inferred from the simulations, because individual resolution elements do not get to see the same random numbers any more when computations have already started to slightly diverge due to the more subtle reasons mentioned earlier.
@@
The first term in equation (6) represents the intrinsic scatter that comes from the astrophysical processes themselves. The origin of this intrinsic scatter can be broadly divided into two main processes – hierarchical merging on one side, and BH accretion and associated feedback on the other side. The galaxies and their central BHs experience a series of mergers across the cosmological evolution. Since the galaxies do not strictly lie on the scaling relation when they merge, the merger remnant realizes scatter for the galaxies. On the other hand, although the existence of a co-evolution of BHs and their host galaxies may regulate galaxy evolution, making galaxies evolve towards the $M_{\rm BH}-M_{\star}$ relation and reducing the scatter, the BH accretion in itself can also evolve BHs away from the relation and introduce scatter in the scaling relation. The degree to which this happens depends on the physics model for galaxy evolution, and in particular on the feedback for BH growth and star formation. When feedback self-regulates the BH growth, the combination of accretion and feedback can also reduce scatter. Thus, while accretion is a driver of scatter, feedback can in fact act in the opposite direction by reducing it through self-regulation, particularly in accretion-dominated regimes. Aside from these physical processes, another important source for scatter in the $M_{\rm BH}-M_{\star}$ relation can lie in variations of the BH seed mass, especially when these seeds are comparatively massive so that substantial galaxy growth is needed before all memory of the seed variations is lost.
To summarize this section, we can consider a simple model to describe the origin of the variances of the BH mass in the $M_{\rm BH}-M_{\star}$ relation:
$\sigma_{\rm BH}^{2}$ $=\underbrace{\sigma_{\rm BH\,seed}^{2}+\sigma_{\rm Acc\&FB}^{2}+\sigma_{\rm HM}^{2}}_{\rm physical}$ (7)
$+\underbrace{\sigma_{\rm SM,\,BH}^{2}+\sigma_{\rm Numerics,\,BH}^{2}+\beta^{2}\sigma_{\rm SM,\,\star}^{2}+\beta^{2}\sigma_{\rm Numerics,\,\star}^{2}}_{\rm numerical}$
$+\sigma_{\rm oths}^{2}+\sigma_{\rm cross}^{2}.$
Here,
  • $\sigma_{\rm BH\,seed}^{2}$ is the variance coming from the variation of BH seed mass;
  • $\sigma_{\rm Acc\&FB}^{2}$ is the variance due to the variation of BH accretion and feedback processes;
  • $\sigma_{\rm HM}^{2}$ is the variance introduced by variations of the merging history of different halos;
  • $\sigma_{\rm SM,\,BH}^{2}$ is the variance from stochastic modelling of BHs in simulations;
  • $\sigma_{\rm Numerics,\,BH}^{2}$ is the variance from non-linear chaotic behaviour of the cosmological simulations;
  • $\beta^{2}\sigma_{\rm SM,\,\star}^{2}+\beta^{2}\sigma_{\rm Numerics,\,\star}^{2}$ are the corresponding terms for the modelling of the stellar mass, which indirectly enter the inferred variance of the BH mass in the $M_{\rm BH}-M_{\star}$ relation;
  • $\sigma_{\rm oths}^{2}$ is the variance from other processes we do not consider explicitly in this model;
  • and $\sigma_{\rm cross}^{2}$ finally captures cross-covariances between terms.
In this framework, the first three contributions are physical in origin, whereas the subsequent four are numerical. For simplicity, we ignore cross terms in the following analysis.
@@ 3.2 Numerical experiments
To better characterize different sources of scatter in the $M_{\rm BH}-M_{\star}$ relation, we carry out different numerical experiments, using the TNG model as a fiducial test case. We begin by estimating the scatter from numerical variations and stochastic modelling. To this end we run five TNG box simulations with a box size of $25\,h^{-1}{\rm cMpc}$, using identical initial condition but different numbers of compute nodes. The latter is a means to introduce implicit numerical variations because the AREPO code is only binary invariant in its results if the number of compute nodes/cores is kept unchanged for a repeated simulation. Varying this will change the domain decomposition, and thus the exact order in which gravitational forces and hydrodynamical fluxes are added up, both of which lead to floating point round-off variations. In addition, the random numbers that will be applied to individual gaseous resolution elements for the stochastic modelling of star formation will be completely different when the number of cores for parallel execution is changed.
For the five results we then perform an object-by-object subhalo matching and calculate the variance and standard deviation for the sample of matched subhalos. To match subhaloes between simulations, we adopt a straightforward strategy based on spatial proximity and subhalo mass. Since all simulations are initialized with identical conditions and differ only in the number of computational nodes, the large-scale structures evolve nearly identically. For each central galaxy, we search for nearest counterparts within 30 ckpc in three-dimensional space and require their stellar masses to agree within 10%. We restrict the matching to central galaxies (excluding satellites) to ensure robustness. With this method, we achieve a matching fraction of approximately 60–70% across the full central galaxy population, increasing to  80% for the massive galaxies that are the primary focus of this study. For each matched set, we calculate the standard deviation of the quantity of interest (e.g. BH mass) among the five simulations. This provides an estimate of the numerical variation for that particular subhalo. To obtain a representative scatter for a given stellar mass, we group the matched subhaloes into stellar mass bins and report the median of the individual standard deviations within each bin. Since the initial conditions are identical, the variation in galaxy properties for the matched subhalos can be attributed to numerical variations and stochastic modelling. Note that Genel et al. (2019) reported that the variations in galaxy properties measured in TNG box simulations through a similar method is not sensitive to the employed numerical resolution.
@@
The left panel of Figure 5 shows the standard deviation of the BH mass in our five TNG simulation with identical initial condition, as mentioned above. The galaxies are also fulfilling the selection criteria mentioned in Section 2.2.1. We find that the variations of the BH mass roughly follow a Gaussian distribution, except that there are also some data points far away from zero. Previous work (e.g., Genel et al., 2019) has shown that the stochastic numerical scatter between simulations with identical initial conditions is well described by a Gaussian distribution, and that the rare extreme deviations may be typically caused by technical issues such as matching failures. Including these outliers would substantially overestimate the variance of the BH mass and obscure the underlying scatter. Since our aim is to quantify the typical numerical fluctuations rather than the full statistical spread including rare anomalies, we exclude data points outside a $3\sigma$ region in the analysis. As shown in the right panel of Figure 5, the distribution of the data is well preserved when these outliers are removed. Figure 6 compares the time evolution of the measured variance of the BH mass before and after cleaning of the data as mentioned above. When outliers are removed, the variance shows a well behaved behaviour with a value that is roughly constant across $z=3$ to $z=0$, which is consistent with the findings in Genel et al. (2019). We will thus adopt this data cleaning strategy for estimating the size of the scatter due to numerical noise, denoted $\sigma^{2}_{\rm BH}$ in equation (6). Subtracting this from the measured full scatter $\hat{\epsilon}^{2}$, we can then also obtain an estimate of the intrinsic physical scatter $\epsilon^{2}$, which can be viewed as being composed of two main sources, hierarchical merging and gas accretion.
@@
We have additionally run five further TNG box simulations using a similar treatment as mentioned above, but turning the BH accretion off. For these simulations, variations due to BH accretion are not introduced by construction, such that the scatter in these simulations is only caused by hierarchical merging and numerical variations. Assuming that the scatter from different sources is approximately uncorrelated and additive, we can estimate the scatter just from BH accretion by subtracting the corresponding result from the full scatter obtained earlier. Here we for simplicity ignore the stellar mass dependence of the scatter, so this is evidently only a coarse estimate. By using this approach, we can separately estimate the contributions to the intrinsic scatter from hierarchical merging and gas accretion.
Table 2: Description of the numerical experiments designed to isolate different components of the scatter model (Equation (6)).
simulations descriptions
TNGFid a 25 Mpc/$h$ box simulation with $2\times 256^{3}$ resolution elements, using the fiducial TNG model
TNGSeries five TNGFid simulations with identical setup but different number of nodes
NoAcc The same as TNGFid simulations but turning of BH accretion
NoAccSeries five NoAcc simulations with identical setup but different number of nodes
Table 3: Measured variances of BH mass from the different numerical experiments, corresponding to the terms in the scatter model (Equation (6)).
simulations descriptions
TNGFid $\sigma_{\rm Acc\&FB}^{2}+\sigma_{\rm HM}^{2}+\sigma_{\rm SM,\,BH}^{2}+\sigma_{\rm Numerics,\,BH}^{2}+\beta^{2}\sigma_{\rm SM,\,\star}^{2}+\beta^{2}\sigma_{\rm Numerics,\,\star}^{2}+\sigma_{\rm oths}^{2}$
TNGSeries $\sigma_{\rm SM,\,BH}^{2}+\sigma_{\rm Numerics,\,BH}^{2}+\beta^{2}\sigma_{\rm SM,\,\star}^{2}+\beta^{2}\sigma_{\rm Numerics,\,\star}^{2}+\sigma_{\rm oths}^{2}$
NoAcc $\sigma_{\rm HM}^{2}+\sigma_{\rm Numerics,\,BH}^{2}+\beta^{2}\sigma_{\rm SM,\,\star}^{2}+\beta^{2}\sigma_{\rm Numerics,\,\star}^{2}+\sigma_{\rm oths}^{2}$
NoAccSeries $\sigma_{\rm Numerics,\,BH}^{2}+\beta^{2}\sigma_{\rm SM,\,\star}^{2}+\beta^{2}\sigma_{\rm Numerics,\,\star}^{2}+\sigma_{\rm oths}^{2}$
As described above, each experiment isolates a different subset of the variance terms in our scatter model (Equation 6). We summarize the design of the experiments and the corresponding variances in Tables 2 and 3. Linear combinations of these results then allow us to estimate the individual contributions of the components $\sigma_{\rm Acc\&FB}^{2}$, $\sigma_{\rm HM}^{2}$, $\sigma_{\rm SM,\,BH}^{2}$, $\sigma_{\rm Numerics,\,BH}^{2}$, $\sigma_{\rm SM,\,\star}^{2}$, and $\sigma_{\rm Numerics,\,\star}^{2}$ in the model.
@@
Figure 6: The time evolution of the variance of the BH mass in matched halos of five identical TNG simulation with and without data cleaning. The figure indicates that after removing data outliers the scatter in the BH mass from numerical origins shows a time-independent behaviour.